OpenStudy (anonymous):

If a resistor draws 1.2 * 10^-3 A of current from a 12 V battery, then what is the value of the resistor? A. 10 ohms B. 100 ohms C. 1.0 kilo-ohms D. 10 kilo-ohms

OpenStudy (michele_laino):

using the Ohm's law, we have: I=12/(0.0012)=...amperes

OpenStudy (michele_laino):

oops..no amperes, Ohms

OpenStudy (anonymous):

10,000?

OpenStudy (michele_laino):

I rewrite: R=12/(0.0012)=...ohms

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

what does that convert to? divide by 1000? and the solution is 10 kilo-ohms? choice D?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay! thanks! :D onto the next!:)

OpenStudy (michele_laino):

:) ok!